Today is Aug 28th and I'll assign one second of homework. On Aug 29th, I'll assign two seconds of homework. On Aug 30, four seconds; on Aug 31, 8 seconds and so on. I'll continue doubling your homework for 30 school days until Oct 11th. From Oct 11th until the end of the year you will have no homework...sounds good, right? How much homework will you end up with. Show your work.
Thanks,
Nick

Answer by stanbon(75887)

1st day: 1^0=1 second
2nd day: 2^1=2 seconds
3rd day: 2^2=4 seconds
4th day: 2^3=8 seconds
....
20th day: 2^19 seconds
30th day: 2^29= 536,870,912 seconds
If you mean "how much time will you have to spend on homework
on the 30th day" the answer is 536,870,912 seconds
Cheers,
Stan H.